# algorithmic

### v4

@@ 1 Introduction
  • The core objective of DP is to evaluate the distinguishability between processed and original data. Yet, a significant hurdle is the difficulty of providing adversarial worst-case proofs, especially when comparing two distinct datasets. This difficulty presents a substantial barrier to validating the efficacy of DP methods in practical scenarios.
@@
Recently, Xiao & Devadas (2023) introduces a novel definition of privacy known as Probably Approximately Correct (PAC) Privacy, representing a significant evolution in privacy-preserving methodologies. This definition diverges from traditional Differential Privacy (DP) by focusing on the difficulty of reconstructing data using any measure function, thereby offering broader applicability beyond the adversarial worst-case scenarios considered by DP. PAC privacy’s strength lies in its model-agnostic nature: it does not depend on an in-depth understanding of a model’s internal workings or its training process. This attribute makes it exceptionally versatile, enabling its application across a wide array of models and contexts with relative ease. Furthermore, leveraging PAC privacy comes with a framework that autonomously determines the minimal noise addition necessary for effective data protection (Xiao & Devadas, 2023). This framework simplifies the process of achieving privacy-preserving data handling, making it more accessible for widespread use in the field of data science and beyond.
@@ 3.2 Conditional Diffusion Models
Where $Z$ is a normalizing constant. According to unconditional reverse process,
@@
And then we take the logarithm, it could be written as
@@
where $g$ is the gradient of classifier
$g=\nabla_{x_{t}}\log p_{\phi}(y|x_{t})|_{x=\mu}$. The investigation has led to the conclusion that the conditional transition operator can be closely estimated using a Gaussian. And the Gaussian resembles the unconditional transition operator, with the distinction that its mean is adjusted by the product of the covariance matrix, $\Sigma$, and the vector $g$. This methodology allows for the generation of high-quality, targeted synthetic data.
3.3 PAC Privacy
Definition 3.3.1. ($(\delta,\rho,D)$ PAC Privacy (Xiao & Devadas, 2023)). For a data processing mechanism $\mathcal{M}$, given some data distribution $D$, a measure function $\rho(.,.)$, and a finite set $X^{*}$, we say $\mathcal{M}$ satisfies $(\delta,\rho,D)$-PAC Privacy if the following experiment is impossible: A user generates data $X$ from distribution $D$ and sends $\mathcal{M(X)}$ to an adversary. The adversary who knows $D$ and $\mathcal{M}$ is asked to return an estimation $\tilde{X}\in X^{*}$ on X such that with probability at least $(1-\delta)$, $\rho(\tilde{X},X)=1$.
Definition above depicts the reconstruction hardness for the attackers to recover the private data distribution $\mathcal{M(X)}$. With a lower bound probability $(1-\delta)$, such that the measure function $\rho(.,.)$ is not able to distinguish the recovery data from the original data. However, the limitation of the naive definition is that the prior distribution of the public dataset is unknown, resulting to the failure of adversarial inferences.
Hence, the author introduces mutual information to measure the mutual dependence of two variables,
where $\mathcal{D}_{KL}$ denotes the KL divergence. Utilizing the mutual information, the Definition can be rewritten as $\mathcal{M}$ satisfies $(\Delta_{KL}\delta,\rho,D)$ PAC Advantage Privacy where
$\Delta_{KL}\delta=D_{KL}(1_{\delta}||1_{\delta_{o}^{\rho}})\leq MI(X;M(X)).$ (6)
Therefore, the goal for the rest of this paper, following Theorem above, is to ascertain or establish limits for the mutual information $MI(X;M(X))$ with a high level of certainty.
@@ 4 Methods
We introduce the PAC Privacy Preserving Diffusion Model (P3DM), which aims to safeguard privacy for specific attributes, building upon the foundation set by DPGEN (Chen et al., 2022). DPGEN introduces a Randomized Response (RR) mechanism, outlined as follows:
$Pr[H(\tilde{x_{i}})=x_{i}^{r}]=\left\{\begin{aligned} &\frac{e^{\epsilon}}{e^{\epsilon}+k-1},&&\omega=x_{i}\\ &\frac{1}{e^{\epsilon}+k-1},&&\omega=x_{i}^{\prime}\in X\setminus x_{i}\end{aligned}\right.$
With the method, $\tilde{x_{i}}$ is designed to point to one of its $k$ nearest neighbors with certain probability $Pr[H(\tilde{x_{i}})=x_{i}^{r}]$, which is designed to enforce $\epsilon$-differential privacy by privatizing the recovery direction $d=(\tilde{x_{i}}-x_{i}^{r})/\sigma^{2}_{i}$. Following perturbation, DPGEN learns an energy score function:
@@
Subsequently, images are synthesized via the Langevin MCMC process.
While DPGEN asserts compliance with stringent $\epsilon$-differential privacy, recent findings Dockhorn et al. (2022) indicate that DPGEN is, in fact, data-dependent. The RR mechanism $M(d)$ in DPGEN holds validity only within the perturbed dataset. Specifically, if an element $z$ belongs to the output set $O$ but not to the perturbed dataset $d$, then $Pr[M(d)=O]=0$, which contravenes the differential privacy definition.
@@ 4.1 Conditional Private Langevin Sampling
where $k$ is the gradient scale that can be tuned according to the performance of the model, $\Sigma_{\theta}(x_{t-1})$ is the covariance matrix from reverse process Eqn. 3. In Eqn. 8, the term $\nabla_{x_{t-1}}\log c_{\theta}(y_{n}|x_{t-1})$ directs the Langevin sampling process toward a specific class label $y_{n}$, which is sampled prior to the inference. This private guidance during the inference phase ensures that the synthesized images are protected from privacy breaches related to designated attributes.
In essence, private classifier guidance can be interpreted as a method of deliberately altering certain generated trajectories. To illustrate, consider an image that is autonomously generated to depict Celebrity A donning eyewear, thereby reflecting the known fact that Celebrity A habitually wears glasses. However, when sampling is conducted with the application of private classifier guidance, the attributes are subject to variation. Consequently, there is a possibility that Celebrity A may be depicted without eyewear in the resulting image. This alteration effectively safeguards the attribute of ’glasses’ from being a consistent feature in the generated representation of Celebrity A. As a result, the images generated under this guidance have a higher privacy extent when compared to those produced by the original DPGEN.
Algorithm 1 Private Conditional Guidance Langevin Dynamics Sampling
0:  class labels $y=\{y_{1},y_{2},...y_{n}\}$ from one of balanced dataset attributes; gradient scale k; $\{\delta_{i}\}_{i=1}^{L}$, $\epsilon$, $T$; pretrained scorenet model $s_{\theta}$;pretrained classifier model on noisy images $c_{\theta}$
@@
7:        $x_{t}$leftarrow $x_{t-1}+\frac{\alpha_{i}}{2}s_{\theta}(x_{t-1},\delta_{i})+\sqrt{\alpha_{i}}z_{t}+k\Sigma_{\theta}(x_{t-1})\nabla_{x_{t-1}}\log c_{\theta}(y_{n}|x_{t-1})$

### v5

@@ 1 Introduction
  • Training modern deep learning models with differential privacy is notoriously difficult, as evidenced by  Tramer & Boneh (2021); He et al. (2023); Ding et al. (2024). Recent progress relies heavily on the use of large pre-trained models (Li et al., 2022; Golatkar et al., 2022; Yu et al., 2022; Li et al., 2022) or extensive public training data (De et al., 2022). We believe that such a paradigm is unlikely to be a holy grail solution, as it does not provide a rigorous privacy guarantee but is more heuristic in nature. We refer the reader to the recent criticism against this paradigm (Tramèr et al., 2024).
@@
Recently, Xiao & Devadas (2023) introduces a novel definition of privacy known as Probably Approximately Correct (PAC) Privacy, representing a significant evolution in privacy-preserving methodologies. PAC Privacy characterizes the information-theoretic hardness to recover sensitive data given arbitrary information disclosure/leakage during/after any processing. Compared with differential privacy, it has the following advantages.
  • Differential privacy requires bounded sensitivity, which cannot be tightly computed for modern machine learning models. Artificial modifications are typically required to decompose algorithms into multiple simpler and analyzable components, e.g., gradient clipping in DP-SGD. The exact privacy analysis of DP can be, in general, NP-hard. Especially for for deep learning, tight DP analysis are challenging. However, PAC privacy can be applied to any data processing algorithm (as a black-box procedure), where security parameters can be produced with arbitrarily high confidence via Monte-Carlo simulation.
  • On the utility side, the magnitude of (necessary) perturbation required in PAC Privacy is not lower bounded by ${\Theta}(\sqrt{d})$ for a $d$-dimensional release, but could be $O(1)$ (depending on data distribution) for many practical data processing tasks, which is in contrast to the input-independent worst-case information-theoretic lower bound like DP. Therefore, PAC Privacy analysis in many applications can produce much sharpened utility-privacy trade-offs.
@@ 3.1 Differential Privacy
3.2 PAC Privacy
3.2.1 Basic Definition
Definition 3.1.
($(\delta,\rho,D)$ PAC Privacy). As discussed in Xiao & Devadas (2023), for a data processing mechanism $\mathcal{M}$, given some data distribution $D$, a measure function $\rho(.,.)$, and a finite set $X^{*}$, we say $\mathcal{M}$ satisfies $(\delta,\rho,D)$-PAC Privacy if the following experiment is impossible: A user generates data $X$ from distribution $D$ and sends $\mathcal{M(X)}$ to an adversary. The adversary who knows $D$ and $\mathcal{M}$ is asked to return an estimation $\tilde{X}\in X^{*}$ on X such that with posterior success probability at least $(1-\delta)$, $\rho(\tilde{X},X)=1$.
Definition 3.2.
($(\Delta_{f}\delta,\rho,\mathsf{D})$ PAC Advantage Privacy) Equivalantly, As discussed in Xiao & Devadas (2023), $\mathcal{M}$ could be defined as $(\Delta_{f}\delta,\rho,\mathsf{D})$ PAC Advantage Privacy if the posterior advantage measured in $f$-divergence $\mathcal{D}_{f}$ satisfies
$\Delta_{f}\delta=\mathcal{D}_{f}(\bm{1}_{\delta}\|\bm{1}_{\delta^{\rho}_{o}})=\delta^{\rho}_{o}f(\frac{\delta}{\delta^{\rho}_{o}})+(1-\delta^{\rho}_{o})f(\frac{1-\delta}{1-\delta^{\rho}_{o}}),$ (2)
where $(1-\delta^{\rho}_{o})$ represents the optimal prior success rate of recovering $X$.
$1-\delta^{\rho}_{o}=\sup_{\tilde{X}\in\mathcal{X}^{*}}\Pr_{X\sim\mathsf{D}}\big{(}\rho(\tilde{X},X)=1\big{)},$ (3)
and $\bm{1}_{\delta}$ and $\bm{1}_{\delta^{\rho}_{o}}$ represent two Bernoulli distributions of parameters $\delta$ and $\delta^{\rho}_{o}$, respectively.
The definition above depicts the reconstruction hardness for the attackers to recover the private data distribution $\mathcal{M(X)}$. With a lower bound probability $(1-\delta)$, the measure function $\rho(.,.)$ cannot distinguish the recovery data from the original data. However, the limitation of the naive definition is that the prior distribution of the public dataset is unknown, resulting in the failure of adversarial inferences.
Definition 3.3.
(Mutual Information). For two random variables $x$ and $w$ in some joint distribution, the mutual information $MI(x;w)$ is defined as
where $\mathcal{D}_{KL}$ denotes the KL divergence.
Theorem 3.1.
As discussed in Xiao & Devadas (2023), for any selected $f$-divergence $\mathcal{D}_{f}$, a mechanism $\mathcal{M}:\mathcal{X}^{*}\to\mathcal{Y}$ satisfies $(\Delta_{f}\delta,\rho,\mathsf{D})$ PAC Advantage Privacy if
$\Delta_{f}\delta=\mathcal{D}_{f}\big{(}\bm{1}_{\delta}\|\bm{1}_{\delta^{\rho}_{o}}\big{)}\leq\inf_{\mathsf{P}_{W}}\mathcal{D}_{f}\big{(}\mathsf{P}_{X,\mathcal{M}(X)}\|\mathsf{P}_{X}\otimes\mathsf{P}_{W}\big{)}.$ (5)
In particular, when we select $\mathcal{D}_{f}$ to be the KL-divergence and $\mathcal{P}_{W}=\mathcal{P}_{\mathcal{M}(X)}$, $\mathcal{M}$ satisfies $(\Delta_{KL}\delta,\rho,\mathsf{D})$ PAC Advantage Privacy where
$\Delta_{KL}\delta=\mathcal{D}_{KL}(\bm{1}_{\delta}\|\bm{1}_{\delta^{\rho}_{o}})\leq MI\big{(}X;\mathcal{M}(X)\big{)}.$ (6)
Proof.
See Appendix A. ∎
In summary, $\mathcal{D}_{f}\big{(}\bm{1}_{\delta}\|\bm{1}_{\delta^{\rho}_{o}}\big{)}$ quantifies the divergence between optimal a priori and posterior reconstruction, effectively measuring the difficulty of inference. A higher value of $\mathcal{D}_{f}\big{(}\bm{1}_{\delta}\|\bm{1}_{\delta^{\rho}_{o}}\big{)}$ signifies greater privacy leakage. Moreover, since $MI\big{(}X;\mathcal{M}(X)\big{)}$ provides an upper bound for $\mathcal{D}_{f}$, a lower value of $MI\big{(}X;\mathcal{M}(X)\big{)}$ indicates stronger privacy protection. Thus, theorem 3.1 establishes a general method for linking the difficulty of arbitrary inference to the well-known concept of mutual information. With theorem 3.1, the goal of PAC privacy is explicit: determining the bound $MI\big{(}X;\mathcal{M}(X)\big{)}$ with high confidence.
3.2.2 Noise Determination and Simulated Privacy Guarantee with High Confidence
The natural idea to achieve information leakage control is perturbation: when $MI\big{(}X;\mathcal{M}(X)\big{)}$ is not small enough to produce satisfied PAC Privacy, we may add additional Gaussian noise $\bm{B}$, to produce smaller $MI\big{(}X;\mathcal{M}(X)+\bm{B}\big{)}$.
Theorem 3.2 ((Xiao & Devadas, 2023)).
When the mutual information $MI(X;M(X))$ is insufficient to ensure PAC privacy, additional Gaussian noise $B\sim\mathcal{N}(0,\Sigma_{B})$ can be introduced to yield a reduced mutual information $MI(X;M(X)+\bm{B})$, such that $MI(X;M(X)+\bm{B})$ satisfies
$\mathsf{MI}(X;\mathcal{M}(X)+\bm{B})\leq\frac{1}{2}\cdot\log\text{det}\big{(}\bm{I}_{d}+\Sigma_{\mathcal{M}(X)}\cdot\Sigma^{-1}_{\bm{B}}\big{)}.$ (7)
In particular, let the eigenvalues of $\Sigma_{\mathcal{M}(\mathbf{X})}$ be $(\lambda_{1},\ldots,\lambda_{d})$, then there exists some $\Sigma_{B}$ such that $\mathbb{E}[\|\bm{B}\|_{2}^{2}]=(\sum_{j=1}^{d}\sqrt{\lambda_{j}})^{2}$, and $\text{MI}(\mathbf{X};\mathcal{M}(\mathbf{X})+\bm{B})\leq 1/2$.
Proof.
See Appendix B. ∎
Therefore, we have a simple upper bound on the mutual information after perturbation which only requires the knowledge of the covariance of $\mathcal{M}(X)$. Another important and appealing property is that the noise calibrated to ensure the target mutual information bound is not explicitly dependent on the output dimensionality $d$ but instead on the square root sum of eigenvalues of $\Sigma_{\mathcal{M}(X)}$. When $\mathcal{M}(X)$ is largely distributed in a $p$-rank subspace of $\mathbb{R}^{d}$, the above theorem suggests that a noise of scale $O(\sqrt{p})$ is needed. Depending on data distribution, if $p$ is a constant, then the noise can be as low as $O(1)$. This is different from DP where the expected $l_{2}$ norm of noise is in a scale of $\Theta(\sqrt{d})$ given constant $L_{2}$-norm sensitivity.
Based on Theorem 3.2, we can use an automatic protocol Algorithm 1 to determine $\Sigma_{B}$ and produce an upper bound such that $\text{MI}(\mathbf{X};\mathcal{M}(\mathbf{X})+\bm{B})\leq(\nu+\beta)$ with high confidence, where $\nu$ and $\beta$ are positive parameters selected as explained below. After sufficiently many simulations, the following theorem ensures that we can obtain accurate enough estimation with arbitrarily high probability.
Theorem 3.3 (Xiao & Devadas (2023)).
Assume that $\mathcal{M}(\mathbf{X})\in\mathbb{R}^{d}$ and $\|\mathcal{M}(\mathbf{X})\|_{2}\leq r$ for some constant $r$ uniformly for any $\mathbf{X}$, and apply Algorithm 1 to obtain the Gaussian noise covariance $\Sigma_{B}$ for a specified mutual information bound $v+\beta$. $v$ and $\beta$ can be chosen independently, and $c$ is a security parameter. Then, there exists a fixed and universal constant $\kappa$ such that one can ensure $\text{MI}(\mathbf{X};\mathcal{M}(\mathbf{X})+\bm{B})\leq\nu+\beta$ with confidence at least $(1-\gamma)$ once the selections of $c$, $m$ and $\gamma$ satisfy,
$c\geq\kappa r\left(\max\left\{\sqrt{\frac{d+\log(4/\gamma)}{m}},\frac{d+\log(4/\gamma)}{m}\right\}+\sqrt{\frac{d\log(4/\gamma)}{m}}\right).$
Another way to interpret this noise determination is that a smaller $\mathbb{E}||B||_{2}$ implies a smaller covariance matrix $\Sigma_{B}$ and lesser noise addition on $M(X)$, which indicates $X$ and $M(X)$ have less mutual information. Therefore, less noise addition signifies a higher privacy protection of the model $M$.
1:  for $k=1,2,\ldots,m$ do
2:     sample images $y^{(k)}=M(X^{(k)})$.
3:  end for
4:  Calculate empirical mean $\hat{\mu}=\frac{1}{m}\sum_{k=1}^{m}y^{(k)}$ and the empirical covariance estimation $\hat{\Sigma}=\frac{1}{m}\sum_{k=1}^{m}(y^{(k)}-\hat{\mu})(y^{(k)}-\hat{\mu})^{T}$.
5:  Apply singular value decomposition (SVD) on $\hat{\Sigma}$ and obtain the decomposition as $\hat{\Sigma}=\hat{U}\hat{\Lambda}\hat{U}^{T}$, where $\hat{\Lambda}$ is the diagonal matrix of eigenvalues $\lambda_{1}\geq\lambda_{2}\geq\ldots\geq\lambda_{d}$.
6:  Determine the maximal index $j_{0}=\arg\max_{j}\lambda_{j}$ for those $\lambda_{j}>c$.
7:  if $\min_{1\leq j\leq j_{0},1\leq d}(\lambda_{j}-\hat{\lambda}_{j})>r\sqrt{d/c+2c}$ then then
8:     for $j=1,2,\ldots,d$ do
9:        Determine the $j$-th element of a diagonal matrix $A_{B}$ as
10:     end for
11:     Determine the Gaussian noise covariance as $\Sigma_{B}=\hat{U}A_{B}\hat{U}^{T}$.
12:  else
13:     Determine the Gaussian noise covariance as $\Sigma_{B}=\left(\sum_{j=1}^{d}\lambda_{j}+dc/(2\nu)\right)\cdot I_{d}$.
14:  end if
15:  Return Gaussian covariance matrix $\Sigma_{B}$.
@@ 3.3 Conditional Diffusion Models
Where $Z$ is a normalizing constant. According to unconditional reverse process, which predicts timestep $x_{t}$ from $x_{t-1}$ leveraging Gaussian distribution, we have
@@
$\log p_{\theta}(x_{t}|x_{t+1})=-\frac{1}{2}(x_{t}-\mu)^{T}\Sigma^{-1}(x_{t}-\mu)+C$ (9)
If we assume that $\log_{\phi}p(y|x_{t})$ has low curvature compared to $\Sigma^{-1}$, then we can approximate $\log p_{\phi}(y|x_{t})$ using a Taylor expansion around $x_{t}=\mu$ as
$\begin{split}\log p_{\phi}(y|x_{t})&\approx\log p_{\phi}(y|x_{t})|_{x_{t}=\mu}+(x_{t}-\mu)\nabla_{x_{t}}\log p_{\phi}(y|x_{t})|_{x_{t}=\mu}\\ &=(x_{t}-\mu)g+C_{1}\end{split}$ (10)
where $g$ is the gradient of classifier $g=\nabla_{x_{t}}\log p_{\phi}(y|x_{t})|_{x=\mu}$ and $C_{1}$ is the constant.
Therefore, combing Eqn. 9 and 10 gives us
@@
The investigation has led to the conclusion that the conditional transition operator can be closely estimated using a Gaussian. And the Gaussian resembles the unconditional transition operator, with the distinction that its mean is adjusted by the product of the covariance matrix, $\Sigma$, and the vector $g$. This methodology allows for the generation of high-quality, targeted synthetic data.
@@ 4 Methods
We introduce the PAC Privacy Preserving Diffusion Model (P3DM), which aims to safeguard privacy for specific attributes. Our method is inspired by DPGEN (Chen et al., 2022). While DPGEN asserts compliance with stringent $\epsilon$-differential privacy, recent findings from Dockhorn et al. (2022) indicate that DPGEN is, in fact, data-dependent. The RR mechanism $M(d)$ in DPGEN holds validity only within the perturbed dataset. Specifically, if an element $z$ belongs to the output set $O$ but not to the perturbed dataset $d$, then $Pr[M(d)=O]=0$, which contravenes the differential privacy definition. Different from DPGEN, which turns out to have no formal privacy guarantees, our method satisfies PAC privacy.
Algorithm 2 PAC-Private Adapted Randomized Response Algorithm
0:  a sensitive dataset $\{x_{i}:i=1,2,...,m\}_{i=1}^{m}$; sample number $k$
1:  Sampling $k$ image candidates from $\{x_{i}\}_{i=1}^{m}$, with the sampling probability for each image Eqn. 13, to construct set $\mathrm{X}$leftarrow $\{x_{j}:max(x_{i}-x_{j})/\sigma_{j}\leq\beta,j\in[m]\}$
2:  Privatize images in $X$ with RR from Eqn. 12 and obtain $H(x_{i})$
3:  $\tilde{x}_{i}=H(x_{i})+B$; $B\sim\mathcal{N}(0,\Sigma_{B})$
4:  Return $\tilde{x}_{i}$
Given a sensitive dataset $\{x_{i}:i=1,2,...,m\}_{i=1}^{m}$, we first sample an image by $\tilde{x}_{i}\leftarrow\mathcal{N}(x_{i},\sigma^{2}I)$. Next, we utilizes the random response method as follows:
$Pr[H(\tilde{x}_{i})=w]=\left\{\begin{aligned} &\frac{e^{\epsilon}}{e^{\epsilon}+k-1},&&w=x_{i}\\ &\frac{1}{e^{\epsilon}+k-1},&&w=x_{i}^{\prime}\in X\setminus x_{i}\end{aligned}\right.$ (12)
In the equation, $\mathrm{X}=\{x_{j}:max(\tilde{x}_{i}-x_{j})/\sigma_{j}\leq\beta,j\in[m]\}$ (where “max” is over the dimensions of $\tilde{x}_{i}-x_{j}$), $|\mathrm{X}|=k\geq 2$, where the hyperparameter $k$ denotes the number of selected candidates from $m$ training images, and the sampling probability for each image is given by
$p(x_{i})=\exp(-d_{\infty}(x_{i},\tilde{x}_{i},\sigma_{i}))/\sum_{a=1}^{m}\exp(-d_{\infty}(x_{i},\tilde{x}_{i},\sigma_{a}))$ (13)
In other words, the mechanism $H(\cdot)$ consists of 2 steps:
  1. 1.
    Sampling $k$ image candidates from $\{x_{i}\}_{i=1}^{m}$, with the sampling probability for each image from Eqn. 13, to construct set $X$.
  2. 2.
    Privatize images in $X$ with RR from Eqn. 12.
After applying $H(\cdot)$, we then add Gaussian noise $B$ to the image to get $\tilde{x}_{i}\leftarrow H(x_{i})+B$ for achieving PAC privacy. With these methods, $\tilde{x}_{i}$ outputs one of its $k$ nearest neighbors with certain probability $Pr[H(\tilde{x}_{i})=x_{i}^{r}]$, as in Chen et al. (2022), giving us the privatized the recovery direction $d_{i}^{r}=(\tilde{x}_{i}-x_{i}^{r})/\sigma^{2}_{i}$.
Following perturbation and privatization, we learn an energy function $q_{\theta}(\tilde{x})$ by optimizing the following loss function:
@@
After getting the optimal $\nabla_{x}\log q_{\theta}(\tilde{x})=(\tilde{x}-x^{r})/\sigma^{2}$ from Eqn. 14, we can then synthesize images using the Langevin MCMC sampler with the optimal output $\nabla_{x}\log q_{\theta}(\tilde{x})$ as:
$x_{t}\leftarrow x_{t-1}+\frac{\alpha_{i}}{2}q_{\theta}(x_{t-1})+\sqrt{\alpha_{i}}z_{t},t=0,1,2,...T.$ (15)
where $\alpha_{i}$ denotes the step size.
@@ 4.1 Conditional Private Langevin Sampling
where $k$ is the gradient scale that can be tuned according to the performance of the model, $\Sigma_{\theta}(x_{t-1})$ is the covariance matrix from reverse process Eqn. 8. In Eqn. 16, the term $\nabla_{x_{t-1}}\log c_{\theta}(y_{n}|x_{t-1})$ directs the Langevin sampling process toward a specific class label $y_{n}$, which is sampled prior to the inference.
This private guidance during the inference phase ensures that the synthesized images are protected from privacy breaches related to designated attributes. At its core, this method involves intentionally modifying certain generated trajectories by randomly perturbing the image label $y_{n}$, which is then used by a pretrained classifier to guide the image generation process. This strategy is aimed at diverting certain image attributes that we wish to protect. For instance, consider a scenario where an original dataset image depicting Celebrity A wearing eyewear, a known trait of the celebrity. If our goal is to privatize the attribute of “wearing glasses,” private classifier guidance can effectively achieve this. When sampling occurs under the influence of this guidance, the attributes are likely to vary, creating a chance that the resulting synthesized image based on Celebrity A might be rendered without glasses. This modification effectively protects the attribute of ’glasses’ from being a consistent element in the generated depictions of Celebrity A. Consequently, the images generated with this approach offer a higher degree of privacy compared to those produced by the original DPGEN method.
Algorithm 3 PAC-Private Conditional Guidance Langevin Dynamics Sampling
0:  class labels $y=\{y_{1},y_{2},...y_{n}\}$ from one of balanced dataset attributes; gradient scale k; $\{\delta_{i}\}_{i=1}^{L}$, $\epsilon$, $T$; optimal output $q_{\theta}$ from Eqn. 14;pretrained classifier model on noisy images $c_{\theta}$
@@
7:        $x_{t}$leftarrow $x_{t-1}+\frac{\alpha_{i}}{2}q_{\theta}(x_{t-1},\delta_{i})+\sqrt{\alpha_{i}}z_{t}+k\Sigma_{\theta}(x_{t-1})\nabla_{x_{t-1}}\log c_{\theta}(y_{n}|x_{t-1})$